# Ionisation Calculations using Classical Molecular Dynamics

### v1

@@ Abstract
To correctly evaluate the bulk properties of WDM in atomistic simulations, large systems are often required. For example, many particles are needed to characterise transport behaviour in the hydrodynamic limit [32, 18, 33]. Therefore, to extend OCP and TCP methods to partially ionised materials, while retaining the computational efficiency necessary to access substantial system sizes, a chemical picture may be required. In this framework each ionic charge state is considered as a distinct species with appropriate inter-particle potentials, such that the equilibrium charge state distribution must be known to compute both dynamic and static equilibrium properties. Importantly, the distribution will depend on the model being used, so must be computed in a self-consistent framework. Ebeling and Militzer attempted such a calculation for a hydrogen plasma wavepacket model by performing simulations to minimise the internal energy as a function of ionisation state at constant entropy, however their approach assumes an ideal entropy term and is therefore limited to weak coupling [21]. In the same work, transitions between bound and free populations were also considered, for which the equilibrium ionisation state becomes an important benchmark. A chemical picture MD approach could also provide a useful reference for TCP modelling strategies that incorporate bound states dynamically such as those discussed in refs. [29, 34].
@@ II Hydrogen models
$\mathcal{H}^{(\rm{MD})}=\sum_{k}^{N_{\rm{i}}}\frac{\mathbf{p}^{\rm{i}}_{k}}{2m_{\rm{i}}}+\sum_{l}^{N_{\rm{n}}}\frac{\mathbf{p}^{\rm{n}}_{l}}{2m_{\rm{n}}}+\mathcal{U}^{(\rm{MD})},$ (3)
where $m_{\rm{i}}$ is the ion mass, and the neutral mass $m_{\rm{n}}=m_{\rm{i}}+m_{\rm{e}}$ is the sum of electron and ion masses. The momentum of the $k^{\rm{th}}$ ion is $\mathbf{p}^{\rm{i}}_{k}$ and the $l^{\rm{th}}$ neutral is $\mathbf{p}^{\rm{n}}_{l}$. The potential energy function $\mathcal{U}^{(\rm{MD})}$ determines the interactions of the system and depends on the configuration of the particles - modifications to this function serve as the central object for the Thermodynamic Integration computation given in section III.1. In the following subsections, the potentials used to model the particle interactions are introduced.
@@ II.3 Short Range Repulsion model
is applied in combination with the BSC model, that acts solely between neutral pairs. In equation (18), $\epsilon$ is the overall constant determining the strength of the interaction and is set by equating $U_{\rm{\rm{SRR}}}^{\rm{nn}}(r_{*})=(3/2)k_{B}T$, to give an effective radius $r_{*}$. The total energy due to SRR interactions is
@@
where the sum runs over all pairs of neutral particles. The potentials discussed thus far have been implemented as a new pairstyle into a the LAMMPS codebase [55], in which all subsequent simulations are performed. The MD potential is given by
@@ III Free Energy Minimisation Framework
In any model which distinguishes between bound and free electrons, the ionisation state is key quantity of interest. In hydrogenic systems, the ionisation state takes a particularly simple form as the ratio of the number of ions to the total number of electrons $\bar{z}=N_{\rm{i}}/N$. A system in thermodynamic equilibrium and held at constant density $n=N/V$ and temperature $T$ will exhibit a minima in the Helmholtz free energy:
@@
Equation (21) is valid under the constraint of charge neutrality, $N_{e}=N_{i}$, and the reduced free energy per particle is $f=F/Nk_{B}T$. The preferred ionisation state may therefore be found by computing the free energy across a set of artificial ionisation states, and performing a minimisation procedure. These considerations apply to any chemical picture of hydrogen, and are readily generalisable to heavier elements [37, 38, 56], where different charge states will be present. To our knowledge, this approach has not been employed in MD applications where, given a set of interparticle potentials, the effect of strong coupling can be treated exactly. The methodology to compute the free energy is detailed in the following section.
@@ III.2 Ideal Free Energy
$f_{n}^{\rm{id}}=\bar{x}\log\left[\frac{1}{2}\bar{x}n\Lambda_{n}\right]-\bar{x}\frac{I}{k_{B}T}-\bar{x},$ (26)
@@ IV.1 Free Energy Minimisation of Bound State Coulomb Model
Figure 3 shows the corresponding free energy curves for the BSC model, where all sampled ionisation states are included. The ideal case and reference systems are also plotted, which are parameterised functions and are readily minimised. A finite set of data are available to minimise the neutral free energy, and therefore a quadratic curve is fit to the lowest 3 data points to extract the minima. The error mainly depends upon the resolution in ionisation state, which we estimate to be half the interval over which the quadratic curve is fitted. When including the OCP interactions, the collective binding energy between the ions and free electron background lowers the free energy at higher ionisation states. Therefore, the ionisation state that minimises the free energy increases. Only ion screening is present and the inclusion of an electronic response would cause further increase, as there are additional negative free energy contributions due to screening [61]. The attractive coulomb interactions increasing the ionisation state is a form of Ionisation Potential Depression (IPD), and corresponds to a lowering of the energy difference between the bound state and the continuum. At $\bar{z}=0$, the OCP and ideal systems have the same free energy, given that no ions are present.
@@ IV.2 Including Short Range Repulsion Effects
Hard sphere potentials have been used many times within the chemical picture literature as computational devices [4, 35, 36, 40, 37, 43, 42, 41, 38, 40]. The charge-neutral interaction in this approach simply restricts the volume accesible to charged species’ which is an analytically tractable mechanism and has therefore found prevalent use. For example, Saumon and Chabrier found that, at high densities, the free energy of their model was lower when fully neutral compared to when fully ionised [35]. Given that the system was expected to be fully pressure ionised, an additional hard sphere potential was introduced between neutral species of radius $r=2\rm{a}_{0}$. Although the model presented here does not include the same framework of internal atomic states, or interatomic interactions, we still observe the same phenomenon. Along similar lines, Potekhin introduced an additional unbinding free energy to account for the inner-core screening effect of overlapping atomic orbitals [43]. Furthermore, in many models in the chemical picture, neutral species’ interaction effects are modelled directly as hard spheres [4, 40, 41].
Given these motivations, and as a demonstration of the technique, an additional SRR potential (as described in section II.3) of radius $r_{*}=2$ has been introduced into the neutral-neutral interaction to study its effect on the predicted ionisation state. We note that such a potential could also be used to model Pauli exclusion effects between bound electrons [53, 54] - these are already treated exactly in the ideal free electron gas but are not present in bound-bound or bound-free interactions. A third integration is required evaluate the influence of the harder inter-neutral core on the ionisation state. A characteristic integration curve is given in figure 4 where all curves are suppressed for small $\lambda_{3}$ by the factor of $(\lambda_{3})^{11}$ in the derivative of the potential energy function. The free energies for this condition are displayed in figure 5. The increased atomic repulsion drives the free energy at lower ionisation states up, causing an increase in the minima. The free energy remains below the OCP free energy for higher ionisation states (not shown in figure) due to the attractive neutral-background contributions.
@@ IV.3 Ionisation State across density
The minimisation results from the simulations performed with both the BSC model and SRR model are compiled in figure 6. The ionisation state of the ideal hydrogen plasma decreases with increasing density. This is because the density of accessible free particle states decreases and it becomes energetically favourable to reduce the effective number of species. Including the OCP interactions turns on long-range charged-particle interactions and increases the ionisation state, as discussed in section IV.1. As $r_{s}$ decreases, the coulomb coupling parameter is larger, which means the negative excess free energy contributions are greater, and the system exhibits larger binding between the ions and electron background, encouraging ionisation. At the highest densities, $r_{s}\leq 1$, the free electron degeneracy contributions are dominant and the ionisation state still tends to zero. Common models for the OCP interactions are the Debye-Hückel and ion-sphere models which, when inserted into the minimisation procedure, give $\bar{z}_{\rm{DH}}$ and $\bar{z}_{\rm{IS}}$ respectively. The models are defined in appendix D. When compared with the OCP reference calculation, both models overpredict the ionisation state. The Debye-Hückel model is correct in the weak coupling limit, and therefore good agreenent is found for $r_{s}\gtrsim 5$. The ion-sphere model is more appropriate in the strongly-coupled (high density) regime, where it outperforms the Debye-Hückel prediction when compared against the more accurate OCP prediction. Generally, the overprediction for both models is severe, considering that the magnitude of deviation is similar to the difference between ideal and OCP curves. The deviation is directly related to both models overestimating the magnitude of the excess internal energy, which can be seen in in figure 7.
We now discuss the effect of introducing neutral interactions, starting with the BSC model introduced in II.1. For $r_{s}\gtrsim 4$, the coulomb interactions of the neutrals have a negligible effect and the deviation in ionisation from the ideal case is purely driven by OCP interactions in the plasma. As the density increases, the effect of the neutral interactions cause a small increase the ionisation state. This is an effect of the ion-neutral and neutral-neutral repulsion effects. At the two highest densities considered: $r_{s}=0.5$ and $r_{s}=0.75$, and for high neutral fractions, the neutral-ion mixture separated into a dense ball of neutral particles surrounded by ions, hence breaking the isotropy of the fluid. This phase separation is another indication that the inter-neutral potential is too weak and these points are omitted from figure 6. We note that at higher ionisation states, this effect is not present, and allows the BSC system to be used as an appropriate reference system when performing integration calculations with the SRR potential.
@@ V Conclusion
The results presented from the one component models provide stimulus for improvement. For example, free electron interaction effects can be included either in effective ion-ion potentials [6, 62, 8], or in two-component models [25]. Free electrons provide additional screening and therefore would further lower the free energy at higher ionisation states [61], causing the equilibrium ionisation state to increase. Meanwhile, multispecies systems with excited states and multiple charge states can be included within the minimisation framework.
Generally, the development of models in the chemical picture could be particularly useful in developing computational frameworks for Ionisation Potential Depression (IPD) [2], given that known models faulter at strong coupling [48], and that IPD is non-trivial to extract from first-principles computations [63, 64, 65, 66]. In the results presented, an IPD effect is present, and dominated by ion-ion coulomb interactions at lower densities. Furthermore, MD provides a promising route to ionic transport coefficients, provided reliable interparticle potentials are supplied.
@@ Appendix B Parameterisation of OCP data
where the boundary condition $f^{\rm{OCP}}(0)=0$ has been applied. There are two small discrepancies between lines 2 and 4 in equation (40) and equivalent terms in equation (16) in ref. [69], which we assume are due to misprints in the latter.
@@ Appendix D Ionisation models
Both DH and IS models apply to the OCP subsystem and don’t include neutral interactions. To estimate the effect of neutrals, the Carnahan-Stirling equation of state may be utilised [70]. Neutrals interacting through the SRR potential with the characteristic lengthscale $r_{*}$ may, in a simple approximation, be modelled as hard spheres of radius $R=r_{*}/2$. In the model, the excess free energy due to hard-sphere interactions is

### v2

@@ Abstract
To correctly evaluate the bulk properties of WDM in atomistic simulations, large systems are often required. For example, many particles are needed to characterise transport behaviour in the hydrodynamic limit [32, 18, 33, 34]. Therefore, to extend OCP and TCP methods to partially ionised materials, while retaining the computational efficiency necessary to access substantial system sizes, a chemical picture may be required. In this framework each ionic charge state is considered as a distinct species with appropriate inter-particle potentials, such that the equilibrium charge state distribution must be known to compute both dynamic and static equilibrium properties. Importantly, the distribution will depend on the model being used, so must be computed in a self-consistent framework. Ebeling and Militzer attempted such a calculation for a hydrogen plasma wavepacket model by performing simulations to minimise the internal energy as a function of ionisation state at constant entropy, however their approach assumes an ideal entropy term and is therefore limited to weak coupling [21]. In the same work, transitions between bound and free populations were also considered, for which the equilibrium ionisation state becomes an important benchmark. A chemical picture MD approach could also provide a useful reference for TCP modelling strategies that incorporate bound states dynamically such as those discussed in refs. [29, 35].
@@ II Hydrogen models
$\mathcal{H}^{(\rm{MD})}=\sum_{k}^{N_{\rm{i}}}\frac{{(\mathbf{p}^{\rm{i}}_{k})^{2}}}{2m_{\rm{i}}}+\sum_{l}^{N_{\rm{n}}}\frac{{(\mathbf{p}^{\rm{n}}_{l})^{2}}}{2m_{\rm{n}}}+\mathcal{U}^{(\rm{MD})},$ (3)
where $m_{\rm{i}}$ is the ion mass, and the neutral mass $m_{\rm{n}}=m_{\rm{i}}+m_{\rm{e}}$ is the sum of electron and ion masses. The momentum of the $k^{\rm{th}}$ ion is $\mathbf{p}^{\rm{i}}_{k}$ and the $l^{\rm{th}}$ neutral is $\mathbf{p}^{\rm{n}}_{l}$. The potential energy function $\mathcal{U}^{(\rm{MD})}$ determines the interactions of the system and depends on the configuration of the particles – modifications to this function serve as the central object for the Thermodynamic Integration computation given in section III.1. In the following subsections, the potentials used to model the particle interactions are introduced.
@@ II.3 Short Range Repulsion model
is applied in combination with the BSC model, that acts solely between neutral pairs. In equation (18), $\epsilon$ is the overall constant determining the strength of the interaction and is set by equating $U_{\rm{\rm{SRR}}}^{\rm{nn}}(2r_{*})=(3/2)k_{B}T$, to give an effective radius $r_{*}$. The total energy due to SRR interactions is
@@
where the sum runs over all pairs of neutral particles. The potentials discussed thus far have been implemented as a new pairstyle into the LAMMPS codebase [56], in which all subsequent simulations are performed. The MD potential is given by
@@ III Free Energy Minimisation Framework
In any model which distinguishes between bound and free electrons, the ionisation state is key quantity of interest. In hydrogenic systems, the ionisation state takes a particularly simple form as the ratio of the number of ions to the total number of electrons $\bar{z}=N_{\rm{i}}/N$. A system in thermodynamic equilibrium and held at constant density $n=N/V$ and temperature $T$ will exhibit a minimum in the Helmholtz free energy:
@@
Equation (21) is valid under the constraint of charge neutrality and $f=F/Nk_{B}T$ is the reduced free energy per particle. The equilibrium ionisation state may therefore be found by computing the free energy across a set of artificial ionisation states and performing a minimisation procedure. These considerations apply to any chemical picture of hydrogen, and are readily generalisable to heavier elements [38, 39, 57], where different charge states will be present. To our knowledge, this approach has not been employed in MD applications where, given a set of interparticle potentials, the effect of strong coupling can be treated exactly. The methodology to compute the free energy is detailed in the following section.
@@ III.2 Ideal Free Energy
$f_{n}^{\rm{id}}=\bar{x}\log\left[\frac{1}{2}\bar{x}n\Lambda_{n}^{3}\right]-\bar{x}\frac{I}{k_{B}T}-\bar{x},$ (26)
@@ IV.1 Free Energy Minimisation of Bound State Coulomb Model
Figure 3 shows the corresponding free energy curves for the BSC model, where all sampled ionisation states are included. The ideal case and reference systems are also plotted, which are parameterised functions and are readily minimised. A finite set of data are available to minimise the neutral free energy, and therefore a quadratic curve is fit to the lowest 3 data points to extract the minima. The error mainly depends upon the resolution in ionisation state, which we estimate to be half the interval over which the quadratic curve is fitted. When including the OCP interactions, the collective binding energy between the ions and free electron background lowers the free energy at higher ionisation states. Therefore, the ionisation state that minimises the free energy increases. Only ion screening is present and the inclusion of an electronic response would cause further increase, as there are additional negative free energy contributions due to screening [62]. The attractive coulomb interactions increasing the ionisation state is a form of ionisation potential depression, and corresponds to a lowering of the energy difference between the bound state and the continuum, a connection which is asserted in appendix A. At $\bar{z}=0$, the OCP and ideal systems have the same free energy, given that no ions are present.
@@ IV.2 Including Short Range Repulsion Effects
Hard sphere potentials have been used many times within the chemical picture literature as computational devices [4, 36, 37, 41, 38, 44, 43, 42, 39, 41]. The charge-neutral interaction in this approach simply restricts the volume accesible to charged species’ which is an analytically tractable mechanism and has therefore found prevalent use. For example, Saumon and Chabrier found that, at high densities, the free energy of their model was lower when fully neutral compared to when fully ionised [36]. Given that the system was expected to be fully pressure ionised, an additional hard sphere potential was introduced between neutral species with radius $2\rm{a}_{0}$. Although the model presented here does not include the same framework of internal atomic states, or interatomic interactions, we still observe the same phenomenon. Along similar lines, Potekhin introduced an additional unbinding free energy to account for the inner-core screening effect of overlapping atomic orbitals [44]. Furthermore, in many models in the chemical picture, neutral species’ interaction effects are modelled directly as hard spheres [4, 41, 42].
Given these motivations, and as a demonstration of the technique, an additional SRR potential (as described in section II.3) with effective radius $r_{*}=1$ has been introduced into the neutral-neutral interaction to study its effect on the predicted ionisation state. We note that such a potential could also be used to model Pauli exclusion effects between bound electrons [54, 55] – these are already treated exactly in the ideal free electron gas but are not present in bound-bound or bound-free interactions. The appropriate form of the repulsive potential is unknown, but could possibly be deduced through comparison with high fidelity data, sourced from experiment or first-principles computations. Alternatively, a scheme based on the pairwise Pauli interactions from wave packet molecular dynamics could be devised [19, 25]. A third integration is required evaluate the influence of the harder inter-neutral core on the ionisation state. A characteristic integration curve is given in figure 4 where all curves are suppressed for small $\lambda_{3}$ by the factor of $(\lambda_{3})^{11}$ in the derivative of the potential energy function. The free energies for this condition are displayed in figure 5. The increased atomic repulsion drives the free energy at lower ionisation states up, causing an increase in the minima. The free energy remains below the OCP free energy for higher ionisation states (not shown in figure) due to the attractive neutral-background contributions.
@@ IV.3 Ionisation State across density
The minimisation results from the simulations performed with both the BSC model and SRR model are compiled in figure 6. The ionisation state of the ideal hydrogen plasma decreases with increasing density.This is because the density of free particle states decreases and it becomes energetically favourable to reduce the effective number of species. Including the OCP interactions turns on long-range charged-particle interactions and increases the ionisation state, as discussed in section IV.1. As $r_{s}$ decreases, the coulomb coupling parameter is larger, which means the negative excess free energy contributions are greater, and the system exhibits larger binding between the ions and electron background, encouraging ionisation. At the highest densities, $r_{s}\leq 1$, the free electron degeneracy contributions are dominant and the ionisation state still tends to zero. Common models for the OCP interactions are the Debye-Hückel and ion-sphere models which, when inserted into the minimisation procedure, give $\bar{z}_{\rm{DH}}$ and $\bar{z}_{\rm{IS}}$ respectively. The models are defined in appendix D. When compared with the OCP reference calculation, both models overpredict the ionisation state. The Debye-Hückel model is correct in the weak coupling limit, and therefore good agreement is found for $r_{s}\gtrsim 5$. The ion-sphere model is more appropriate in the strongly-coupled (high density) regime, where it outperforms the Debye-Hückel prediction when compared against the more accurate OCP prediction. Generally, the overprediction for both models is severe, considering that the magnitude of deviation is similar to the difference between ideal and OCP curves. The deviation is directly related to both models overestimating the magnitude of the excess internal energy, which can be seen in in figure 7.
We now discuss the effect of introducing neutral interactions, starting with the BSC model introduced in II.1. For $r_{s}\gtrsim 4$, the coulomb interactions of the neutrals have a negligible effect and the deviation in ionisation from the ideal case is purely driven by OCP interactions in the plasma. As the density increases, neutral interactions induce a small increase in the ionisation state. This is an effect of the ion-neutral and neutral-neutral repulsion effects. At the two highest densities considered: $r_{s}=0.5$ and $r_{s}=0.75$, and for high neutral fractions, the neutral-ion mixture separated into a dense ball of neutral particles surrounded by ions, hence breaking the isotropy of the fluid. This phase separation is another indication that the inter-neutral potential is too weak and these points are omitted from figure 6. We note that at higher ionisation states, this effect is not present, and allows the BSC system to be used as an appropriate reference system when performing integration calculations with the SRR potential.
@@ V Conclusion
The results presented from the one component models provide stimulus for improvement. For example, free electron interaction effects can be included either in effective ion-ion potentials [6, 63, 8], or in two-component models [25]. Free electrons provide additional screening and therefore would further lower the free energy at higher ionisation states [62], causing the equilibrium ionisation state to increase. Meanwhile, multispecies systems with excited states and multiple charge states can be included within the minimisation framework. On including these developments, and leveraging the presented ionisation state calculation framework, comparison could be made against hydrogen equation of state tables [2].
Generally, the development of models in the chemical picture could be particularly useful in developing computational frameworks for ionisation potential depression (IPD) [2], given that known models falter at strong coupling [49], and that IPD is non-trivial to extract from first-principles computations [64, 65, 66, 67]. In the results presented, an IPD effect is present, and dominated by ion-ion coulomb interactions at lower densities. Furthermore, molecular dynamics provides a promising route to ionic transport coefficients, provided reliable interparticle potentials are supplied.
@@ Appendix B Parameterisation of OCP data
where the boundary condition $f^{\rm{OCP}}(0)=0$ has been applied. There is a small discrepency between line 4 in equation (40) and the equivalent term in ref. [70], which we assume is due to a misprint in the latter, while the remaining terms formally agree.
@@ Appendix D Ionisation models
Both DH and IS models apply to the OCP subsystem and don’t include neutral interactions. To estimate the effect of neutrals, the Carnahan-Stirling equation of state may be utilised [71]. Neutrals interacting through the SRR potential with the effective radius $r_{*}$ may, in a simple approximation, be modelled as hard spheres of radius $R=r_{*}$. In the model, the excess free energy due to hard-sphere interactions is